5(a) A company is discussing a $0.5 million loan with a bank. The interest rate is 12% compounded annually and the repayment period is 5 years. The bank is offering two options for loan repayment: Option A: Payments are to be received in equal installments at the end of each year Option B: Interest is to be received on a yearly basis and the Principal is to be received at the end All loan repayment items are end-of-year payments Which options is better for the company? For Option A, construct a spreadsheet that shows for each period the following: annual payment, interest received, recovered capital, and unrecovered capital. For Option B, construct a spreadsheet that shows for periods 1 to 4, interest payments, and for the 5th (last) period, interest as well as the principal. 5(b) A bond with a face value of $500,000 pays quarterly interest of 1.5% per each period. Twenty interest payments remain before the bond matures. How much would you be willing to pay for this bond today if the next interest payment is due now and you want earn 8% compounded quarterly on your money?
Answer #1

5A)

Option A)

First of all lets calculate Installment amount

Installment amount = Loan/PVIFA(r%,n)

=500000/PVIFA(12%,5 years)

=500000/3.6047

=138705$

Statement showing repayment schedule

Towards
Year Opening balance Installment Interest @ 12% Principal Closing balance
1 500000 138705 60000 78705 421295
2 421295 138705 50555 88150 333145
3 333145 138705 39977 98728 234418
4 234418 138705 28130 110575 123843
5 123843 138705 14861 123844 0
193524 500000

Thus total Interest = 193,524$

Option B)

Statement showing repayment of principal and Interest

Year Interest payments Principal payment Total
1 60000 60000
2 60000 60000
3 60000 60000
4 60000 60000
5 60000 500000 560000
Total 300000 500000

Here Interest is 300,000$

Since Interest payment in option A is less than option B. Option A is preferable

5B)

First of all lets find EAR

EAR = (1+ rate/n)^n-1

(1+8%/4)^4-1

=1.02^4-1

=1.08243-1

=8.243%

Means Quarterly rate = 8.243%/4 = 2.0608%

Statement showing PV of interest payment as well as principle

Period Interest Principal Total payment PVIF @ 2.0608% Present Value
0 7500 7500 1.0000 7500
1 7500 7500 0.9798 7349
2 7500 7500 0.9600 7200
3 7500 7500 0.9406 7055
4 7500 7500 0.9216 6912
5 7500 7500 0.9030 6773
6 7500 7500 0.8848 6636
7 7500 7500 0.8669 6502
8 7500 7500 0.8494 6371
9 7500 7500 0.8323 6242
10 7500 7500 0.8155 6116
11 7500 7500 0.7990 5993
12 7500 7500 0.7829 5872
13 7500 7500 0.7671 5753
14 7500 7500 0.7516 5637
15 7500 7500 0.7364 5523
16 7500 7500 0.7215 5412
17 7500 7500 0.7070 5302
18 7500 7500 0.6927 5195
19 7500 500000 507500 0.6787 344442
463783

Thus price of bond that should be paid = 463783$